### v2

# Resolving Gauge Ambiguities of the Berry Connection in Non-Hermitian Systems

Abstract
Non-Hermitian systems display spectral and topological phenomena absent in Hermitian physics; yet, their geometric characterization can be hindered by an intrinsic ambiguity rooted in the eigenspace of non-Hermitian Hamiltonians, which becomes especially pronounced in the pure quantum regime. Because left and right eigenvectors are not related by conjugation, their norms are not fixed, giving rise to a biorthogonal ${\rm GL}(N,{\mathbb{C}})$ gauge freedom. Consequently, the standard Berry connection admits four inequivalent definitions depending on how left and right eigenvectors are paired, giving rise to distinct Berry phases and generally complex-valued holonomies. Here we show that these ambiguities and the emergence of complex phases are fully resolved by introducing a covariant-derivative formalism built from the metric tensor of the Hilbert space of the underlying non-Hermitian Hamiltonian. The resulting uniquely defined Berry connection remains real-valued under an arbitrary ${\rm GL}(N,{\mathbb{C}})$ frame change, and transforms as an affine gauge potential, while reducing to the conventional Berry (or Wilczek–Zee) connection in the Hermitian limit. This establishes an unambiguous and gauge-consistent geometric framework for Berry phases, non-Abelian holonomies, and topological invariants in quantum systems described by non-Hermitian Hamiltonians.
I Introduction
Berry connections, Berry phases, holonomies, and the topology encoded in Bloch bands play a central role in modern quantum physics, underpinning phenomena ranging from charge pumping and the quantum Hall effect to topological insulators and superconductors [8, 40, 43, 44, 19, 34, 31]. Their non-Hermitian generalizations have recently attracted considerable attention due to the emergence of unique spectral and topological structures that cannot occur in Hermitian systems [4, 17, 24, 7].
Non-Hermitian Hamiltonians naturally emerge in a wide variety of experimental platforms, including photonic systems with engineered gain and loss [32, 14, 31], mechanical and electrical metamaterials with active or nonreciprocal elements [16, 35], acoustic lattices [48, 21], cold-atom systems with controlled dissipation [42, 4], circuit-QED and microwave platforms [33], as well as synthetic quantum matter  [31, 7] to name a few. In these systems, the nontrivial phenomena such as non-Hermitian skin effects [47, 25, 27] and non-Bloch band topology [47, 9, 30] demonstrate that their geometric properties encode unique features not encountered in Hermitian counterparts.
A defining feature of the non-Hermitian Hamiltonians is the biorthogonality of their eigenvectors: right and left eigenstates are independent, and need not be related by Hermitian conjugation [28, 10]. As a consequence, each eigenmode retains an intrinsic ‘normalization’ gauge freedom, namely,
$|\psi^{R}_{n}\rangle\rightarrow c_{n}(\boldsymbol{\lambda})|\psi^{R}_{n}\rangle,\qquad\langle\psi^{L}_{n}|\rightarrow c_{n}(\boldsymbol{\lambda})^{-1}\langle\psi^{L}_{n}|,$ (1)
with $c_{n}(\boldsymbol{\lambda})\in\mathbb{C}\setminus{0}$, such that the transformed eigenstates remain biorthonormal $\langle\psi^{L}_{n}|\psi^{R}_{n}\rangle=1$. This residual gauge freedom is far richer than the unitary $U(N)$ gauge of Hermitian systems and is mathematically encoded by the non-compact general linear group ${\rm GL}(N,\mathbb{C})$.
While this freedom leaves spectral properties invariant, it directly affects geometric quantities built from derivatives of the eigenstates. For example, for a single-band with in general complex-valued energy $E_{n}$, by pairing its left and right eigenvectors, the Berry connection admits four nonequivalent forms:
${\cal A}^{\alpha\beta}_{\mu,n}=i\langle\psi^{\alpha}_{n}|\partial_{\mu}\psi^{\beta}_{n}\rangle,\qquad\alpha,\beta\in{L,R},$ (2)
whose transformations under $c_{n}(\boldsymbol{\lambda})$ differ because left and right vectors transform oppositely [15, 37, 41]. As a result, these four connections are generically not gauge equivalent [45]. Moreover, for any parameter-dependent scaling with $|c_{n}|\neq 1$, all ${\cal A}^{\alpha\beta}$ acquire different gauge-dependent imaginary components, rendering the associated Berry phases complex and curvatures non-unique [37]. Specifically, under such rescaling in Eq. (1) with $c_{n}=r_{n}e^{i\phi_{n}}$, the left-right Berry connection for a single-band transforms as
${\cal A^{\prime}}^{LR}_{\mu,n}(\boldsymbol{\lambda})={\cal A}^{LR}_{\mu,n}(\boldsymbol{\lambda})-\partial_{\mu}\phi_{n}+i\partial_{\mu}\ln r_{n}.$ (3)
Clearly, for an arbitrary real-valued scaling factor $r_{n}\neq 1$, the Berry connection becomes complex-valued even if the original connection ${\cal A}^{LR}$ is purely real. This implies that one can induce arbitrary geometric exponential damping or amplification of an adiabatically evolving state along open-path trajectories merely by rescaling the right eigenframe, regardless of whether the spectrum of the underlying non-Hermitian Hamiltonian is real or not.
While such complex geometric phases pose no conceptual difficulty in classical and semiclassical dissipative systems, where the evolving state can be regarded as a physical field that may naturally attenuate or amplify along a given trajectory in parameter space [38], and where such geometric effects can even be observed and exploited, as recently demonstrated in Refs. [41, 26], the situation is fundamentally different in the quantum domain. When the wave function encodes probability amplitudes of the system’s possible configurations, rather than a physical field, its norm represents the total probability and must therefore remain fixed under any admissible continuous evolution, whether generated by a Hermitian or a non-Hermitian operator [13, 5]. Although it is common to enforce norm preservation through ad hoc renormalization of the quantum state, this procedure raises concerns about the consistency of such an approach in genuine non-Hermitian quantum settings [39]. These difficulties have motivated recent efforts to formulate the real-valued Berry connection for non-Hermitian systems [45, 46]; however, the applicability of such approaches has so far been restricted to systems with degenerate eigenspaces.
Consequently, this naturally raises the question of whether one can define a unique and real-valued Berry connection for finite-dimensional non-Hermitian Hamiltonians that applies uniformly to both non-degenerate and degenerate eigenspaces and remains real under arbitrary ${\rm GL}(N,\mathbb{C})$ frame transformations. Such a connection, which ensures the geometric norm preservation of the adiabatically evolving states, would also lead to a consistent geometric and topological structure on the eigenvector bundle of the quantum non-Hermitian evolution operators. In this work, we construct such a Berry connection and show that it naturally emerges from the metric associated with the deformed non-Hermitian Hamiltonian Hilbert space.
II Metric tensor and covariant derivative in the Hilbert space of non-Hermitian Hamiltonians
Consider dynamics of a state $|\psi\rangle$ evolving under a non-Hermitian Hamiltonian (NHH) $H(\boldsymbol{\lambda})$ defined over the parameter space $\boldsymbol{\lambda}$ through the Schrödinger equation (by setting $\hbar=1)$:
$i\partial_{t}|\psi\rangle=H(\boldsymbol{\lambda})|\psi\rangle.$ (4)
To treat $|\psi\rangle$ as a quantum state describing the probability amplitudes, we require its norm to remain unity. For a NHH this necessitates a positive-definite metric operator $\eta$ on its Hilbert space such that
$\langle\psi|\eta|\psi\rangle=1.$ (5)
To ensure the latter the metric itself must obey the equation $\partial_{t}\eta=i\eta H-iH^{\dagger}\eta$. Evidently, for Hermitian systems $\eta=I$. Because $\eta$ is Hermitian and positive-definite operator, it can be alternatively represented as
$\eta=S^{\dagger}S,$ (6)
for some invertible operator $S$, which is in general time-dependent. The operator $S$ plays the role of the vielbein within the framework of the Einstein elevator [23]. Because of this, one can effectively map the right vectors $|\psi\rangle$ of NHH to a vector
$|\phi^{H}\rangle\equiv S|\psi\rangle,$ (7)
corresponding to, in general, a time-dependent Hermitian Hamiltonian $H^{H}=SHS^{-1}+i(\partial_{t}S)S^{-1}$ [23]. The latter expression coincides with the gauge transformation law for the Hamiltonian viewed as a gauge potential, with the transformation rendering it Hermitian. Indeed, combining Eqs. (5)–(7) one immediately obtains $\langle\phi^{H}|\phi^{H}\rangle=1$, which is a standard definition of the inner product on the Hilbert space of Hermitian Hamiltonians (see also Fig. 1111Note that the Hermitizing map $S$ possesses a unitary gauge freedom and thus is not unique [23]. Indeed, by setting $S\to US$, with some unitary $U$, the metric $\eta$ in Eq. (6) remains invariant. To avoid this ambiguity, one can choose a minimal $S$, whose parameter dependence stems solely from the metric, i.e., $S(\boldsymbol{\lambda})=\sqrt{D(\boldsymbol{\lambda})}W^{\dagger}(\boldsymbol{\lambda})$, such that $\eta(\boldsymbol{\lambda})=WDW^{\dagger}$, with $W\in U(N)$, and $D$ is a positive-definite diagonal matrix. As such, here, we implicitly assume that if there is such a $U$, it is always parameter independent.
Consequently, in an $\eta$-deformed Hilbert space the inner product would vary across the system parameter space $\boldsymbol{\lambda}$, rendering ordinary derivatives insufficient for comparing state vectors at different points. A covariant derivative $D_{\lambda}=\partial_{\lambda}+\Gamma^{\lambda}$ is thus required [29]. The covariant derivative, in comparison to the ordinary derivative, additionally incorporates the metric-compatible connection $\Gamma^{\lambda}$ (not to be confused with the Berry connection) that captures the ‘deformation’ of the Hilbert space and ensures that the derivatives of states are defined consistently on the associated Hilbert space bundle.
Figure 1: Symbolic illustration of the vielbein map $S$, according to Eq. (7), which relates the deformed (e.g., stretched) Hilbert space ${\cal H}_{\rm NH}$ of the non-Hermitian Hamiltonian, equipped with the metric $\eta=S^{\dagger}S$, and characterized by nonequivalent local frames at different points, to the flat Hilbert space ${\cal H}_{\rm Herm}$ associated with the corresponding Hermitian Hamiltonian, which admits a single global frame.
The construction of the covariant derivative is greatly simplified by exploiting the relation between the states of non-Hermitian and Hermitian Hamiltonians in Eq. (7). Since the ordinary derivative is well defined on the flat Hilbert space of $H^{H}$ one can pull back this structure to the $\eta$-deformed Hilbert space of the NHH $H$, giving a metric-compatible covariant derivative of the form [3]
$D_{\lambda}=\partial_{\lambda}+S^{-1}\partial_{\lambda}S,$ (8)
where the $S^{-1}\partial_{\lambda}S=\Gamma^{\lambda}$ term serves as the connection component on the $\eta$-deformed Hilbert space, and which obeys the metric-compatible equation for the connection: $\partial_{\lambda}\eta=\eta\Gamma^{\lambda}+\Gamma^{\lambda\dagger}\eta$  [^2] More specifically, $\partial_{\theta}\langle\psi|\psi\rangle_{\eta}=\langle D_{\theta}\psi|\psi\rangle_{\eta}+\langle\psi|D_{\theta}\psi\rangle_{\eta}$=0, implying $\langle D_{\theta}\psi|\psi\rangle_{\eta}=-\langle\psi|D_{\theta}\psi\rangle_{\eta}$. Importantly, this condition does not enforce parallel transport of the state, determined by the expression $D_{\lambda}|\psi\rangle=0$, but merely guarantees norm preservation of a state when it is transported in $\boldsymbol{\lambda}$-space [22]. This situation is fully analogous to the Hermitian case ($\Gamma\to 0$): although the norm of an eigenstate $|\phi^{H}\rangle$ is fixed throughout parameter space $\partial_{\lambda}\langle\phi^{H}|\phi^{H}\rangle=0$, the state itself generally varies with system parameters, i.e., $\partial_{\lambda}|\phi^{H}\rangle\neq 0$. It is precisely this parametric variation that determines the Berry connection..
Furthermore, this connection is curvature-free, as it is generated by a pure-gauge structure (i.e., the Maurer-Cartan form) $S^{-1}dS$, implying $D_{\mu},D_{\nu}]=0 [$29, 22]. Meaning that the Hilbert space, endowed with the metric $\eta$, is intrinsically flat despite the seeming deformation (see Fig. 1).
III Covariant Berry connection
In the adiabatic limit, geometric properties of the cyclic dynamics, e.g., the accumulated Berry phase, are entirely determined by the geometry of the eigenbundle constructed from the eigenvectors of the non-Hermitian Hamiltonian [38, 41]. This statement remains valid for pseudo-Hermitian systems, including those hosting exceptional points (EPs) [2], as well as for non-Hermitian systems with complex spectra whose eigenvalues do not exhibit EP-induced branching, except scenarios where such a branching does not induce the nonadiabatic transitions during the state evolution [36].
With the covariant derivative in Eq. (8), we can now introduce the covariant Berry connection (CBC) $\mathbb{A}$ on, in general, the finite $N$-dimensional degenerate eigenvector space of the non-Hermitian Hamiltonian:
${\mathbb{A}}^{\lambda}_{mn}=i\langle\psi^{R}_{m}|\eta D_{\lambda}|\psi^{R}_{n}\rangle,$ (9)
or alternatively via left and right eigenvectors as
${\mathbb{A}}^{\lambda}_{mn}=i\langle\psi^{L}_{m}|\partial_{\lambda}|\psi^{R}_{n}\rangle+i\langle\psi^{L}_{m}|\Gamma^{\lambda}|\psi^{R}_{n}\rangle,$ (10)
where the left $\langle\psi^{L}|$ and right $|\psi^{R}\rangle$ eigenvectors of the NHH obey the eigenvalue equation $\langle\psi^{L}_{k}|H=\langle\psi^{L}_{k}|E$ and $H|\psi^{R}_{k}\rangle=E|\psi^{R}_{k}\rangle$, $k=1,\dots,N$, for a given $N$-fold degenerate energy $E$, and we exploited the fact that $\langle\psi^{L}|=\langle\psi^{R}|\eta$.
In the Hermitian limit, for a fixed eigenframe, the CBC can be reduced to the standard Berry connection, since $\eta\to I$ and $D_{\lambda}\to\partial_{\lambda}$. Likewise, when the metric constant, that is $\Gamma^{\lambda}=0$, the CBC is simplified to the usual left-right non-Hermitian Berry connection given in Eq. (2). In this sense, the standard non-Hermitian left–right connection ${\cal A}^{LR}$ is appropriate in the genuine quantum regime only when the Hilbert-space metric is fixed (i.e., $\eta=\rm const$). Moreover, via the similarity transformation in Eq. (7), the CBC can also be written as ${\mathbb{A}}^{\lambda}_{mn}=i\langle\phi^{H}_{m}|\partial_{\lambda}|\phi^{H}_{n}\rangle$, which is the conventional Berry connection for a Hermitian system. Thus the CBC in Eq. (9) is manifestly real-valued.
Now, let $|R\rangle$ ($\langle L|$) denote the matrices consisting of columns (rows) of right (left) degenerate eigenvectors above. One can verify that, for such an $N$-dimensional eigenvector frame of the NHH $H$, an arbitrary frame transformation $T\in{\rm GL}(N,\mathbb{C})$: $|R^{\prime}\rangle=|R\rangle T$ and $\langle L^{\prime}|=T^{-1}\langle L|$ induces the following affine gauge transformation of the CBC:
${\mathbb{A}}^{\prime\lambda}=T^{-1}\tilde{\mathbb{A}}^{\lambda}T+iT^{-1}\partial_{\lambda}T,\quad\tilde{\mathbb{A}}^{\lambda}={\mathbb{A}}^{\lambda}+\Xi^{\lambda}.$ (11)
where
$\Xi^{\lambda}=i\langle R|\eta\left(\Gamma^{\prime\lambda}-\Gamma^{\lambda}\right)|R\rangle,$ (12)
and $\Gamma^{\prime}$ is the connection on the Hilbert space of the NHH spanned by the new eigenvectors $|R^{\prime}\rangle$. Equation (11) show that, in contrast to the standard gauge transformation law of the conventional Berry connection, the CBC acquires an additional affine contribution $\Xi^{\lambda}$, reflecting the frame-dependent geometry of the Hilbert space under non-unitary transformations. This affine connection structure echoes the metric-affine gauge theories of gravity, where affine (or distortion) contributions to the connection naturally emerge from metric deformations under frame transformations [20].
Here, the distortion tensor $\Xi$ is determined by the difference of the metric connections $\Gamma^{\prime\lambda}-\Gamma^{\lambda}$, originating from the non-unitary (positive-definite) part $P$ in the polar decomposition of $T=UP$. In other words, the distortion compensates for the local rescaling of eigenvector norms induced by $P$. In particular, for a purely unitary transformation $T=U$, the metric remains invariant $\eta^{\prime}=\eta$, which makes $(\Gamma^{\prime}-\Gamma)$ a pure gauge  [^3] As noted earlier in the text, for the fixed metric $\eta$, the two connections $\Gamma$ and $\Gamma^{\prime}$ may arise from the two Hermitized maps $S$ and $S^{\prime}=US$, respectively, which differ by a smooth unitary $U$. According to the definition of the connection in Eq. (8), these two connections therefore differ by a pure gauge term, $\Gamma^{\prime}-\Gamma=U^{\dagger}{d}U$. By choosing $U$ to be parameter-independent, this gauge contribution vanishes, ensuring that $\Gamma^{\prime}=\Gamma$., implying that one can always set $\Xi^{\lambda}=0$, and thus $\mathbb{A}_{\lambda}$ transforms as an ordinary $U(N)$ gauge potential.
Crucially, the affine gauge transformation law in Eq. (11) guarantees that the CBC remains real under any ${\rm GL}(N,\mathbb{C})$ frame change. Indeed, it is straightforward to show that Eq. (11) is equivalent to
${\mathbb{A}}^{\prime\lambda}=i(\Phi^{\prime H})^{\dagger}\partial_{\lambda}\Phi^{\prime H},$ (13)
where $\Phi^{\prime H}$ is the unitary matrix formed by the eigenstates $|\phi^{\prime H}_{k}\rangle=S^{\prime}|\psi^{\prime R}_{k}\rangle$ of the associated Hermitian Hamiltonian, thus ensuring that the transformed CBC ${\mathbb{A}}^{\prime\lambda}$ stays real. Here $\Phi^{\prime H}=S^{\prime}S^{-1}\Phi^{H}T$, with $\Phi$ denoting the matrix of Hermitian eigenstates $|\phi^{H}_{k}\rangle$ defined in Eq. (7). From this perspective, the affine form of the CBC transformation in Eq. (11) can be interpreted as arising from a combined left–right action on $\Phi^{H}$. That is, the affine term $\Xi$ originates from the left action $S^{\prime}S^{-1}$, which renormalize the eigenstates after a general ${\rm GL}(N,\mathbb{C})$ frame transformation $T$, and which thus highlights the departure of Eq. (11) from the standard gauge potential transformation under right action only. In this regard, the difference of the metric connections determining the distortion tensor in Eq. (12) can be compactly expressed as
$\Gamma^{\prime\lambda}-\Gamma^{\lambda}=S^{-1}(M^{-1}\partial_{\lambda}M)S,$ (14)
where the operator $M=\Phi^{\prime H}T^{-1}\Phi^{H\dagger}$ encodes the transition between two Hermitizing maps $S^{\prime}=MS$, and in general $M\notin U(N)$  [^4] Indeed, since the metric is uniquely determined by the left eigenvectors as $\eta=L^{\dagger}L$, the transformed metric follows $\eta^{\prime}=L^{\prime\dagger}L^{\prime}=L^{\dagger}(T^{-1})^{\dagger}T^{-1}L$. Given $\eta^{\prime}=S^{\prime\dagger}S^{\prime}$ and $L=\Phi^{H\dagger}S$ [the latter comes from $\Phi^{H}=SR=SL^{-1}$ according to Eq. (6)], the Hermitizing map transforms as $S^{\prime}=\Phi^{\prime H}T^{-1}L=\Phi^{\prime H}T^{-1}\Phi^{H\dagger}S$, from which one gets $M=\Phi^{\prime H}T^{-1}\Phi^{H\dagger}$. .
To recap, enforcing a self-consistent state–norm preservation in the Hamiltonian Hilbert space necessitates a redefinition of the Berry connection and a corresponding modification of its transformation law. The resulting affine gauge transformation rule, given in Eq. (11), faithfully incorporates the change of the underlying metric of the Hilbert space and ensures that the modified Berry connection transforms covariantly under general ${\rm GL}(N,\mathbb{C})$ change of the eigenbasis.
IV Berry curvature
Evidently, the Berry curvature associated with the CBC has exactly the same algebraic structure as in Hermitian systems. In the general differential form and component notation, respectively, it reads [44]
$F={\rm d}{\mathbb{A}}-i\mathbb{A}\wedge\mathbb{A},\quad F_{\nu\mu}=\partial_{\nu}{\mathbb{A}}^{\mu}-\partial_{\mu}{\mathbb{A}}^{\nu}-i[{\mathbb{A}}^{\nu},{\mathbb{A}}^{\mu}],$ (15)
where $\rm d$} denotes the ordinary exterior derivative acting on the connection one-form, and $\wedge$ is the graded wedge product for matrix-valued differential forms [18]. This immediately implies that, given the affine transformation law Eq. (11), the Berry curvature transforms accordingly under a general frame change $T\in{\rm GL}(N,\mathbb{C})$ as
$F^{\prime}=T^{-1}\tilde{F}T,\qquad\tilde{F}={\rm d}\tilde{\mathbb{A}}-i\tilde{\mathbb{A}}\wedge\tilde{\mathbb{A}}$, (16)
with $\tilde{\mathbb{A}}$ given in Eq. (11). Importantly, the affine curvature $\tilde{F}$ represents the genuine gauge-covariant curvature of the eigenbundle associated with norm-preserving state vectors under frame transformations.
Because the affine gauge potential and the associated field strength defined in Eqs. (11) and (16), respectively, are uniquely defined, the corresponding topological invariants, such as Chern numbers, are likewise uniquely defined. This is in contrast to four different definitions of the Chern numbers for Abelian eigenbundles discussed in Ref. [37], which nevertheless give the same topological invariant. Furthermore, all contributions involving the affine distortion $\Xi$ in Eq. (16) are either zero or enter the Chern integrals as globally exact differential forms and therefore automatically vanish, since the distortion tensor solely stems from the curvature-free connections $(\Gamma,\Gamma^{\prime})$ in Eq. (8) [1].
V Example
Let us illustrate the proposed framework with a concrete example. We consider a pseudo-Hermitian Hamiltonian of the form:
$H^{\rm pH}=t\sigma_{z}+ix\sigma_{y}-iy\sigma_{x}+a_{0}I,\quad t,x,y,a_{0}\in{\mathbb{R}},$ (17)
where $\sigma_{i}$ are Pauli matrices. This two-level Hamiltonian can, for instance, effectively describe topological magnons on a honeycomb lattice [11, 6]. The same model was recently analyzed in Ref. [12], in the context of emerging electromagnetic fields in the eigenspace of equivalent pseudo-Hermitian Hamiltonians. Here, by contrast, we show that when the system is treated in the purely quantum regime and analyzed within the present covariant formalism, no synthetic electromagnetic fields emerge in such systems.
The eigenvalues of the Hamiltonian are $E_{\pm}=a_{0}\pm l$, with $l=\sqrt{t^{2}-x^{2}-y^{2}}$, implying that the condition $l>0$ defines an exact pseudo-Hermitian phase where the eigenspectrum is real. Through the hyperbolic parameterization $t=l\cosh\xi$, $x=l\sinh\xi\cos\lambda$, $y=l\sinh\xi\sin\lambda$, the eigenvectors take the form on the reduced 2-dimensional parameter manifold $(\lambda,\xi)$:
$|\psi^{R}_{-}\rangle=\begin{pmatrix}-e^{-i\lambda}\sinh\frac{\xi}{2}\\ \cosh\frac{\xi}{2}\end{pmatrix},\quad|\psi^{L}_{-}\rangle=\begin{pmatrix}e^{-i\lambda}\sinh\frac{\xi}{2}\\ \cosh\frac{\xi}{2}\end{pmatrix}$,
$|\psi^{R}_{+}\rangle=\begin{pmatrix}e^{-i\lambda}\cosh\frac{\xi}{2}\\ -\sinh\frac{\xi}{2}\end{pmatrix},\quad|\psi^{L}_{+}\rangle=\begin{pmatrix}e^{-i\lambda}\cosh\frac{\xi}{2}\\ \sinh\frac{\xi}{2}\end{pmatrix}$. (18)
Following Ref. [12], for simplicity, we focus on a single isolated energy band with $E_{-}$. By fixing $E_{-}=const\neq a_{0}$, and exploiting Eqs. (2) and (V), the corresponding conventional left-right Berry connection reads
$A^{LR}_{-}=i\langle\psi^{L}_{-}|{\rm d}|\psi^{R}_{-}\rangle=-\sinh^{2}\frac{\xi}{2}d\lambda.$ (19)
Now, using the biorthogonal gauge freedom for the eigenvectors in Eq. (V), one may rescale both of them by a normalized factor $N=\sinh^{2}\frac{\xi}{2}+\cosh^{2}\frac{\xi}{2}$  [^5] The normalization factor $N$ is chosen such that $\langle\psi^{\prime R}_{i}|\psi^{\prime R}_{i}\rangle=1$ for $i=\pm$., namely $|\psi^{\prime R}_{i}\rangle=N^{-\frac{1}{2}}|\psi^{R}_{i}\rangle$ and $|\psi^{\prime L}_{i}\rangle=N^{\frac{1}{2}}|\psi^{L}_{i}\rangle$, such that $\langle\psi^{\prime L}_{i}|\psi^{\prime R}_{i}\rangle=\langle\psi^{L}_{i}|\psi^{R}_{i}\rangle=1$ for $i=\pm$. That is, in this particular case the frame transformation is $T=NI_{2}\in{\rm GL}(2,\mathbb{R})$. Because of this, and in accordance with Eq. (3), the conventional left-right Berry connection transforms as
$A^{\prime LR}_{-}=A^{LR}_{-}+\frac{i}{2}(\tanh\xi)d\xi.$ (20)
Evidently, after such a rescaling the left–right Berry connection acquires an additional imaginary term. As a consequence, along a given open-path trajectory the Berry phase develops a dissipative component $\gamma_{d}=\int(\tanh\xi)\mathrm{d}\xi$. This shows that within the conventional non-Hermitian formalism the biorthogonal gauge freedom alone can induce a geometric ‘amplification’ or ‘attenuation’ of a state along open paths in the system parameter space, even though the energy spectrum itself remains entirely real.
Now we consider the CBC and demonstrate how it gauges away such an imaginary contribution. For that purpose, one must first find the connection $\Gamma=S^{-1}dS=\Gamma^{\lambda}d\lambda+\Gamma^{\xi}d\xi$, according to Eqs. (8) and (9). The components of the connection are found as (see Appendix A for details):
$\Gamma^{\lambda}=\begin{pmatrix}i&0\\ 0&0\end{pmatrix},\quad\Gamma^{\xi}=\frac{1}{2}\begin{pmatrix}0&e^{-i\lambda}\\ e^{i\lambda}&0\end{pmatrix}$. (21)
Consequently, the CBC in Eq. (9) reads (see Appendix B):
${\mathbb{A}}_{-}=0.$ (22)
That is, under the wave-function norm–preserving condition, the CBC vanishes identically. As a result, no geometric phase is accumulated when the state undergoes a cyclic evolution. This is in contrast to the conventional left–right Berry connection in Eq. (19), which generally depends on $\xi$.
Crucially, under the rescaling of the eigenvectors by the normalization factor $N$ introduced above, the associated CBC, according to Eq. (11), remains unchanged, i.e.,
${\mathbb{A}}^{\prime}_{-}=0,$ (23)
see Appendix C for details. This demonstrates that the CBC is insensitive to such length-changing transformations of the eigenvectors. Indeed, the affine component of the CBC in Eq. (12), $\Xi=-\frac{i}{2}\tanh\xi d\xi$ (see Appendix C), exactly cancels the arising imaginary contribution in the left-right Berry connection in Eq. (20), thus ensuring norm preservation of the quantum state under adiabatic evolution.
In light of the above, it is also clear that, for this two-level model, under the considered constraint $E_{-}=const$, no Berry curvature is generated within the CBC formalism, i.e., $F=0$ according to Eq. (15), in the entire $(\lambda,\xi)$-space. This is in contrast to the nonzero Berry curvature computed from the left-right Berry connection in Eq. (19)  [^6] Specifically, the Berry curvature computed from the conventional left-right Berry connection in Eq. (19) reads $F=-i(\sinh\xi)\sigma_{y}/2$ [12].. Consequently, in comparison with Ref. [12], no ‘fields’ emerge in the eigenspace of such a pseudo-Hermitian system in the quantum regime. We note, however, that the emerging fields, studied in Ref. [12], might still manifest when the Hamiltonian effectively describes (semi)classical systems, where the norm preservation is not essential.
The findings above can be additionally understood from the perspective of the Hermitizing maps $S$ and $S^{\prime}$ associated with the original and rescaled eigenvectors in Eq. (V). Indeed, as shown in Appendix D, the corresponding eigenvectors of the Hermitian Hamiltonians, $H^{H}=SH^{\rm pH}S^{-1}\equiv l\sigma_{x}$, and $H^{\prime H}=S^{\prime}H^{\rm pH}S^{\prime-1}\equiv l\cosh\xi\sigma_{x}$, namely $|\phi^{H}_{-}\rangle=S|\psi^{R}_{-}\rangle$ and $|\phi^{\prime H}_{-}\rangle=S^{\prime}|\psi^{\prime R}_{-}\rangle$, are equivalent up to an arbitrary phase, with the globally smooth $S$ and $S^{\prime}$. Explicitly, one finds
$|\phi^{\prime H}_{-}\rangle\equiv|\phi^{H}_{-}\rangle=\frac{1}{\sqrt{2}}\bigl[1,1\bigr]^{T}.$ (24)
That is, the eigenvectors associated with the Hermitian Hamiltonian are fixed throughout the entire two-dimensional system parameter space $(\xi,\lambda)$, and the corresponding Berry connections, therefore, vanish identically in this region.
VI Conclusions
In this work, we have introduced a uniquely defined, real-valued Berry connection that characterizes the geometric and topological structure of the eigenspace of non-Hermitian Hamiltonians in the genuine quantum regime. By incorporating the parameter-dependent Hilbert-space metric and associated with it the covariant derivative, the resulting covariant framework properly accounts for the general ${\rm GL}(N,\mathbb{C})$ frame transformations, ensuring that the CBC remains real-valued, and eliminates gauge ambiguities inherent to the conventional biorthogonal Berry connection. This provides a consistent and unambiguous geometric framework for defining both the Abelian and non-Abelian holonomies, curvatures, and topological invariants in quantum systems described by non-Hermitian Hamiltonians. We additionally illustrated the proposed framework using a known two-level pseudo-Hermitian model, demonstrating that the geometric structure encoded in its Hamiltonian eigenspace exhibits a substantial qualitative difference between the classical and quantum regimes. This distinction highlights the central role of quantum-state norm preservation, properly incorporated via the covariant formalism, in defining a physically consistent Berry connection and curvature for non-Hermitian systems.
Acknowledgements.
The author would like to thank Chia-Yi Ju for insightful discussions. The author also acknowledges support from the Air Force Office of Scientific Research (AFOSR) Award No. FA8655-24-1-7376, from the Ministry of Education, Youth and Sports of the Czech Republic Grant OP JAC No. CZ.02.01.01/00/22_008/0004596, and from the Grant Agency of the Czech Republic (Project No. 25-15775S).
Appendix A Calculating metric-compatible connection $\Gamma$ for the Hilbert space of pseudo-Hermitian Hamiltonian in Eq. (17)
Here we elaborate on the explicit forms of the metric-compatible connections $\Gamma$ and $\Gamma^{\prime}$ for the Hilbert space of the pseudo-Hermitian Hamiltonian in Eq. (17) spanned by the original eigenvectors in Eq. (V), and by the eigenvectors rescaled by the real factor $N$ introduced below Eq. (19), respectively.
Let us start with the connection
$\Gamma=S^{-1}\mathrm{d}S=\Gamma^{\lambda}d\lambda+\Gamma^{\xi}d\xi.$ (25)
To construct it explicitly, one first needs to determine the Hermitizing map $S$. This map (up to some redundant unitary gauge) can be obtained using the standard decomposition [3],
$S=\sqrt{D}U^{\dagger},$ (26)
where $D$ and $U$ are diagonal and unitary matrices, respectively, arising from the spectral decomposition of the metric $\eta=S^{\dagger}S$, which is diagonalized as $\eta=UDU^{\dagger}$ (away from the exceptional points). The metric is first attained via the expression
$\eta=\sum\limits_{i=\pm}|\psi^{L}_{i}\rangle\langle\psi^{L}_{i}|.$ (27)
Now combining Eqs. (27) and (V) one finds
$\eta=\begin{pmatrix}\cosh\xi&e^{-i\lambda}\sinh\xi\\ e^{i\lambda}\sinh\xi&\cosh\xi\end{pmatrix}$. (28)
The eigenvalues and eigenvectors of the metric are thus read
$D=\begin{pmatrix}e^{-\xi}&0\\ 0&e^{\xi}\end{pmatrix},\quad U=\frac{1}{\sqrt{2}}\begin{pmatrix}-e^{-i\lambda}&e^{i\lambda}\\ 1&1\end{pmatrix}$. (29)
As a result, the Hermitizing map attains the form (up to an arbitrary U(2) gauge):
$S=\frac{1}{\sqrt{2}}\begin{pmatrix}-e^{-\frac{\xi}{2}+i\lambda}&e^{-\frac{\xi}{2}}\\ e^{\frac{\xi}{2}+i\lambda}&e^{\frac{\xi}{2}}\end{pmatrix}$. (30)
Utilizing Eqs. (25) and (30), one then arrives at Eq. (21).
Appendix B Calculating the covariant Berry connection in Eq. (22)
According to Eq. (9), the CBC corresponding to the eigenenergy $E_{-}$ is defined as
${\mathbb{A}}_{-}=i\langle\psi^{R}_{-}|\eta D_{\lambda}|\psi^{R}_{-}\rangle d\lambda+i\langle\psi^{R}_{-}|\eta D_{\xi}|\psi^{R}_{-}\rangle d\xi.$ (31)
Exploiting Eq. (21), one obtains its components as follows
${\mathbb{A}}^{\lambda}=$ $i\langle\psi^{L}_{-}|\partial_{\lambda}|\psi^{R}_{-}\rangle+i\langle\psi^{L}_{-}|\Gamma^{\lambda}|\psi^{R}_{-}\rangle$ (32)
$=(-\sinh^{2}\frac{\xi}{2}+\sinh^{2}\frac{\xi}{2})d\lambda=0,$
and
${\mathbb{A}}^{\xi}=i\langle\psi^{L}_{-}|\partial_{\xi}|\psi^{R}_{-}\rangle+i\langle\psi^{L}_{-}|\Gamma^{\xi}|\psi^{R}_{-}\rangle=(0+0)d\xi=0.$ (33)
Consequently, one arrives at Eq. (22) with $\mathbb{A}=0$.
Appendix C Transformed covariant Berry connection in Eq. (23)
Let us now explicitly compute the transformed CBC $\mathbb{A}^{\prime}$ using the modified metric-compatible connection on the Hilbert space $\Gamma^{\prime}=S^{\prime-1}\mathrm{d}S^{\prime}$. Combining this expression with (37) and (21), one finds the corresponding components
$\Gamma^{\prime\lambda}=\Gamma^{\lambda}=\begin{pmatrix}i&0\\ 0&0\end{pmatrix},\quad\Gamma^{\prime\xi}=\frac{1}{2}\begin{pmatrix}\tanh\xi&e^{-i\lambda}\\ e^{i\lambda}&\tanh\xi\end{pmatrix}$. (34)
Substituting Eq. (34) into Eq. (9), one immediately obtains $\mathbb{A}^{\prime}=0$.
It is thus evident that the invariance of the vanishing CBC under the $T=NI_{2}\in{\rm GL}(2,\mathbb{R})$ transformation is ensured by the appearance of the affine contribution $\Xi$ in Eq. (11). Explicitly, this ‘distortion’ term, in accordance with Eq. (12), is given by
$\Xi$ $=$ $i\langle\psi^{R}_{-}|\eta(\Gamma-\Gamma^{\prime})|\psi^{R}_{-}\rangle$ (35)
$=$ $-\frac{i}{2}\langle\psi^{R}_{-}|\eta\begin{pmatrix}\tanh\xi&0\\ 0&\tanh\xi\end{pmatrix}|\psi^{R}_{-}\rangle$
$=$ $-\frac{i}{2}\tanh\xi\langle\psi^{R}_{-}|\eta|\psi^{R}_{-}\rangle=-\frac{i}{2}\tanh\xi.$
As a result, this affine term exactly compensates the imaginary component that appears in the conventional left–right Berry connection in Eq. (20), thus ensuring that the norm of the wave function is automatically preserved when the state is adiabatically transported in the system parameter space.
Appendix D Hermitized Hamiltonians and their eigenstates
The fact that $\mathbb{A}=0$ can be alternatively understood from the perspective of Hermitized Hamiltonians and their eigenstates. Since the metric, and hence the Hermitizing map, is time independent, which is always the case in the unbroken pseudo-Hermitian phase of the Hamiltonian, the associated Hermitian Hamiltonian can be written as $H^{H}=SH^{\rm pH}S^{-1}$, with the corresponding eigenvectors given by $|\phi^{H}_{i}\rangle=S|\psi^{R}_{i}\rangle$. Explicitly, one finds
$|\phi^{H}_{-}\rangle=\frac{1}{\sqrt{2}}\begin{pmatrix}1\\ 1\end{pmatrix},\qquad|\phi^{H}_{+}\rangle=\begin{pmatrix}-1\\ 1\end{pmatrix}$. (36)
Similarly, one obtains the Hermitizing map for the rescaled eigenvectors as follows
$S^{\prime}=\sqrt{\cosh\xi}S.$ (37)
That is, the new Hermitizing map is just the ‘rescaled’ map $S$. Consequently, the associated eigenvectors $|\phi^{\prime H}_{i}\rangle$ are equivalent to the eigenvectors $|\phi^{H}_{i}\rangle$ in Eq. (36). As a result, the transformed CBC $\mathbb{A}^{\prime}$, in accordance with Eq. (13), is identically zero.